OpenStudy (anonymous):

athcalculus HELP: use your knowledge of the derivative to compute the limit given below: lim (x+h)^(12/12)-(x)^(12/12)/h h->0 the derivative that is being calculated is dy/dx

OpenStudy (anonymous):

isn't \(\frac{12}{12}=1\)

OpenStudy (anonymous):

h/h = 1, limit of 1 as h -> 0 is 1

OpenStudy (anonymous):

well it's square root

OpenStudy (anonymous):

dy/dx=12?

OpenStudy (anonymous):

\[f \left( x \right)=\sqrt[12]{x ^{12}}= x, f'\left( x \right)=1 \text{ so the limit is }1\]

OpenStudy (anonymous):

question: do you only find the derivative of the first number...ex^

OpenStudy (anonymous):

f(x)

OpenStudy (anonymous):

or the whole numerator?

OpenStudy (anonymous):

the limit presented is the difference quotient of the function, which will give you the derivative or slope. the function resolves to f(x) = x, with domain, such that x >= 0. the derivative will be 1 for all x>0.

OpenStudy (anonymous):

i still don't understand.... :/

OpenStudy (anonymous):

for any problem if it's more than zero it = to 1...?

OpenStudy (anonymous):

i tried to find the derivative of that and i got 12.

OpenStudy (anonymous):